### v1

@@ 2 The Probabilistic Adaptive Screening (PAS) method
Multi-fidelity methods connect LF indicators and an HF loads at different statistical levels (see Figure˜2). To the left of this figure: retaining more information leads to greater accuracy but requires complex solution strategies and more HF samples. To the right: simplifying the system reduces HF samples but risks losing critical information. [VS2025] explained that AS was developed at the cumulative distribution level, but in complex cases this approach may have discarded too much information. It may be easier and more robust to model the joint distribution of paired LF-HF peak values than to model differences between their marginal cumulative distributions. In practice, a single LF value may correspond to multiple HF outcomes. This variability is naturally represented in the joint distribution framework but not in the AS method, which implicitly assumes a one-to-one mapping between LF and HF values. In the Probabilistic Adaptive Screening (PAS) method, we therefore apply our model for the LF-HF relationship at the level of the LF and HF peaks instead. We use a copula model to do this. As mentioned, we included the copula basics in Appendix˜B: general theory, formulations of the copulas used in the PAS method, a description of the parameters in each of these copula, and the relation of these parameters to the dependence structure of the fitted data.
@@ 2. Step 2
    LF Monte-Carlo simulation, unchanged: Perform LF Monte-Carlo Simulations (MCS) for a large number $N$ of wave seeds, each with the same exposure duration $T_{\text{exp}}$. This exposure duration is the duration for which you want to obtain the extreme value of the response in a single sea state. This is generally the duration for which the hydrodynamic response is expected to remain ergodic (roughly 15 min to 3 h, depending on wave ergodicity, speed and course changes, etc.). The total MCS should be significantly longer than $T_{\text{exp}}$ to obtain converged extreme values. See [VSS2023, SVS2023] for an example of the required $N$ for wave crests, green water impact forces and wave-in-deck impact forces. The total MCS duration $T_{\text{tot}}$ follows from $T_{\text{tot}}=NT_{\text{exp}}$.
@@ 12. Step 12
The MPM predicted in Step 10 corresponds to the $q=e^{-1}\approx 0.368$ quantile of the short-term distribution for linear Gaussian processes (following the Rayleigh distribution, see e.g., [Ochi1990]). This means that the probability that the maximum response over the target exposure duration exceeds the MPM is 62.3%. The offshore industry often prefers higher quantiles for design. We can estimate these by replacing $P_{\text{exp}}$ in Equation˜11 with $P_{q}=1-q^{m/n}$ ([OTG14]). However, this necessitates an even longer LF MCS to achieve converged results. More research is needed to evaluate the applicability of the method for higher quantiles.
@@ 2.3 Initial sampling
The outcome of Bayesian or adaptive / sequential sampling methods (such as PAS) can depend on the initial sampling, because the initial data shape the first prediction, which in turn influences subsequent updates. If the initial samples do not sufficiently represent the true variability of the system, the prediction may converge more slowly or toward a biased region. To avoid this, we selected a relative large initial sampling pool of 9 LF samples in Step 5 (so also 9 initial corresponding HF samples in Step 6). These LF samples were selected as follows: the 4 largest LF indicator values in the MCS, plus 5 samples that were selected using uniformly spaced indices over the remaining part of ln($\mathbf{d}^{\text{mcs}}_{L})$ from Step 4.
@@ 4 Method used to validate PAS
In a real design procedure for a maritime structure, HF CFD simulations would need to be performed for new wave events (or a few events) at every iteration in Step 6 to obtain the required HF loads. In the present validation study, however, we circumvent this requirement: rather than running CFD simulations, we extract the corresponding HF load responses directly from a validation dataset for each selected wave event at every iteration. This allows us to focus on validating the statistical framework of the PAS method without mixing it with CFD load prediction accuracy (following the same approach used in [VS2025] for AS). For cases 1 and 2, obtaining identical deterministic wave inputs for the LF and HF variables was straightforward. For cases 3 and 4, this was achieved by combining experimentally measured waves recorded next to the ferry, propagated to the vessel’s centre of gravity, with the linear frequency-domain potential flow database of RWE response functions. This ensured that the wave phasing of the LF RWE time traces closely matched that of the experimental waves. These simulations were previously carried out for [VBS2024], where the full procedure is described in detail.
@@
As explained in the introduction, PAS was based on AS. This AS method was published in [VS2025] (with associated scripts in [VSdata2025]), and fine-tuned in [VS2025prads] (with associated scripts in [VSdata2025PRADS]). PAS was compared to AS in the present publication; an overview of the steps in AS that differ from PAS and some relevant formulations are included in Appendix˜A.
@@
Figure 9: Case 2 - VBM: convergence of one-hour MPM as a function of number of samples (top) and final converged distributions (bottom) from AS and PAS. The copula in the name of the PAS results is the used model in the last iteration.
@@ 5 Results and discussion
These figures show that AS and PAS both provide accurate MPM and distribution predictions for all variations of cases 1 and 2. This was also demonstrated in [VS2025] for AS. However, the results obtained using AS with the USMV acquisition function were slightly more accurate than the present AS with MVPD results for case 2a. $M_{2}$ in Figure˜14 shows that both methods produce MPM results with accuracy well within 5% for these two cases, which is acceptable for design purposes. However, the results for cases 3 and 4 clearly highlight the advantage of PAS over AS. The inclusion of probabilistic copula fitting in PAS significantly reduces the bias introduced by the quality of the screening indicator in AS. The $M_{2}$ values show that PAS is able to predict the target extreme MPM values with an accuracy within 10% for all variations of cases 3 and 4, whereas AS shows deviations of up to 40%. The figure also clearly shows the non-conservative tendency of AS (which is problematic for design load estimation). This trend is not present in the results from PAS; the small deviations observed in the MPM and the distributions can be either conservative or non-conservative. Similar results are observed for $M_{3}$, which highlights that the maximum deviations of the PAS distributions compared to the ground truth are within 18% for all test cases, and those of AS within almost 80%.
@@
When comparing the extreme values from the validation datasets with those obtained using (P)AS, the respective uncertainties of these values must be considered. The distributions in Figures˜7, 9, 11 and 13 also show the U95% MF-GPR uncertainty for AS, and the U95% sample uncertainty based on 20 draws from the fitted copula model for PAS. These uncertainty measures are not directly comparable, but they show that the sample uncertainty from PAS is small around $P_{\text{exp}}$ for all eight case variations (whereas the MF-GPR uncertainty band from AS is larger around $P_{\text{exp}}$). For case 1, the HF validation uncertainty is 0.1 m ($\sim$1% of the true MPM value, see Table˜3 and [VS2025]). The U95% uncertainty of the PAS results is 0.06 m (case 1a) and 0.21 m (case 1b), which corresponds to $\sim$0.6% and $\sim$2% of the true MPM values, respectively. For case 3, the root-mean-square errors (RMSE) of the extreme values in the HF validation data are discussed in detail in [VSS2023], showing that these converge to well below 100 kN for the 30-minute MPM in both test conditions. This roughly translates [^2] Assuming a normal distribution, RMSE is equal to standard deviation error, and U95% $\approx 1.96$ RMSE. to an U95% uncertainty around 200 kN ($\sim$7-15% of the true MPM values). The U95% uncertainty of the PAS results is 96 kN (case 3a) and 292 kN (case 3b), which corresponds to $\sim$7% and $\sim$10% of the true MPM values, respectively. The uncertainty of the extreme values in the HF validation data was not evaluated for cases 2 and 4, but it is expected to be between case 1 and 3 for case 2, and comparable to case 3 for case 4. Based on this, both the experimental and PAS uncertainties are considered to be sufficiently small for all cases to make the validation comparison meaningful.
@@
When accuracy would be the only indicator for the performance of an extreme value prediction method, MCS would always be preferred. The objective of (P)AS is to make accurate predictions in a more efficient way. $M_{1}$ in Figure˜14 shows that both methods reduce the required number of HF samples that need to be evaluated significantly compared to the HF validation data: the number of wave events to evaluate with an HF tool is below 1.7% compared to those required in the MCS for all cases and both AS and PAS. The figure also shows that AS is consistently more efficient than PAS. This can be interesting for relatively simple cases such as case 1 and 2, where AS can provide accurate results, and quicker than PAS. PAS is the more reliable method for accurate predictions across all cases, particularly the complex ones, even though it is slightly less efficiency than AS. PAS needs evaluation of less than 0.3% of the HF events compared to MCS for case 1a,b and 2a,b, and less than 1.7% of the HF events compared to MCS for case 3a,b and 4a,b. The exact number of required events depends on the limits set for the stopping criteria. These were defined per case in the present study; the optimal stopping limits for a new case depend on the designer’s requirements, acceptable risks, considered quantiles, and probably requirements by class societies.

### v2

@@ 2 The Probabilistic Adaptive Screening (PAS) method
Multi-fidelity methods connect LF indicators and an HF loads at different statistical levels (see Figure˜2). To the left of this figure: retaining more information leads to greater accuracy but requires complex solution strategies and more HF samples. To the right: simplifying the system reduces HF samples but risks losing critical information. [VS2025] explained that AS was developed at the cumulative distribution level, but in complex cases this approach may have discarded too much information. It may be easier and more robust to model the joint distribution of paired LF-HF peak values than to model differences between their marginal cumulative distributions. In practice, a single LF value may correspond to multiple HF outcomes. This variability is naturally represented in the joint distribution framework but not in the AS method, which implicitly assumes a one-to-one mapping between LF and HF values. In the Probabilistic Adaptive Screening (PAS) method, we therefore apply our model for the LF-HF relationship at the level of the LF and HF peaks instead. We use a copula model to do this, and include the copula basics in Appendix˜B: general theory, formulations of the copulas used in the PAS method, a description of the parameters in each of these copula, and the relation of these parameters to the dependence structure of the fitted data.
@@ 2. Step 2
    LF Monte-Carlo simulation, unchanged: Perform LF Monte-Carlo Simulations (MCS) for a large number $N$ of wave seeds, each with the same exposure duration $T_{\text{exp}}$. This exposure duration is the duration for which you want to obtain the extreme value of the response in a single sea state. This is generally the duration for which the hydrodynamic response is expected to remain ergodic (roughly 15 min to 3 h, depending on wave ergodicity, speed and course changes, etc.). The total MCS should be significantly longer than $T_{\text{exp}}$ to obtain converged extreme values. See [VSS2023, SVS2023] for an example of the required $N$ for wave crests, green water impact forces, and wave-in-deck impact forces. The total MCS duration $T_{\text{tot}}$ follows from $T_{\text{tot}}=NT_{\text{exp}}$.
@@ 12. Step 12
The MPM predicted in Step 10 corresponds to the $q=e^{-1}\approx 0.368$ quantile of the short-term distribution for linear Gaussian processes (following the Rayleigh distribution, see e.g., [Ochi1990]). This means that the probability that the maximum response over the target exposure duration exceeds the MPM is 62.3%. The offshore industry often prefers higher quantiles for design. We can estimate these by replacing $P_{\text{exp}}$ in Equation˜11 with $P_{q}=1-q^{m/n}$ [OTG14]. However, this necessitates an even longer LF MCS to achieve converged results. More research is needed to evaluate the applicability of the method for higher quantiles.
@@ 2.3 Initial sampling
The outcome of Bayesian or adaptive / sequential sampling methods (such as PAS) can depend on the initial sampling, because the initial data shape the first prediction, which in turn influences subsequent updates. If the initial samples do not sufficiently represent the true variability of the system, the prediction may converge more slowly or toward a biased region. To avoid this, we selected a relatively large initial sampling pool of 9 LF samples in Step 5 (so also 9 initial corresponding HF samples in Step 6). These LF samples were selected as follows: the 4 largest LF indicator values in the MCS, plus 5 samples that were selected using uniformly spaced indices over the remaining part of ln($\mathbf{d}^{\text{mcs}}_{L})$ from Step 4.
@@ 4 Method used to validate PAS
In a real design procedure for a maritime structure, HF CFD simulations would need to be performed for new wave events (or a few events) at every iteration in Step 6 to obtain the required HF loads. In the present validation study, however, we circumvent this requirement: rather than running CFD simulations, we extract the corresponding HF load responses directly from a validation dataset for each selected wave event at every iteration. This allows us to focus on validating the statistical framework of the PAS method without mixing it with CFD load prediction accuracy (following the same approach used in [VS2025] for AS). We have shown before that accurate HF wave event impact loads can be obtained with CFD (see e.g., [BHV2020, VMSHKSG2021]). For cases 1 and 2, obtaining identical deterministic wave inputs for the LF and HF variables was straightforward. For cases 3 and 4, this was achieved by combining experimentally measured waves recorded next to the ferry, propagated to the vessel’s centre of gravity, with the linear frequency-domain potential flow database of RWE response functions. This ensured that the wave phasing of the LF RWE time traces closely matched that of the experimental waves. These simulations were previously carried out for [VBS2024], where the full procedure is described in detail.
@@
As explained in the introduction, PAS was based on AS. This AS method was published in [VS2025] (with associated scripts in [VSdata2025]), and fine-tuned in [VS2025prads] (with associated scripts in [VSdata2025PRADS]). The performance of PAS is compared to that of AS in the present publication. An overview of the steps in AS that differ from PAS and some relevant formulations are included in Appendix˜A. It is important to note that, throughout this study, PAS and AS are evaluated using the same analytical and linear potential-flow LF indicators. For brevity, this is not reflected in the method names; whenever AS is referenced, this indicator choice is implied. This represents a deviation from the approach used for pilot case 3 in [VS2025], where the AS method employed an LF indicator derived from coarse-mesh CFD simulations. That AS-coarse mesh CFD combination yielded accurate distributions of green water impact loads on a containership in the pilot study.
@@
Figure 9: Case 2 - VBM: convergence of one-hour MPM as a function of number of samples (top) and final converged distributions (bottom) from AS and PAS. The copula in the names of the PAS results is the used model in the last iteration.
@@ 5 Results and discussion
These figures show that AS and PAS both provide accurate MPM and distribution predictions for all variations of cases 1 and 2. This was also demonstrated in [VS2025] for AS. However, the results obtained using AS with the USMV acquisition function were slightly more accurate than the present AS with MVPD results for case 2a. $M_{2}$ in Figure˜14 shows that both methods produce MPM results with accuracy well within 5% for these two cases, which is acceptable for design purposes. However, the results for cases 3 and 4 clearly highlight the advantage of PAS over AS combined with linear potential flow. The inclusion of probabilistic copula fitting in PAS significantly reduces the bias introduced by the quality of the screening indicator in AS. The $M_{2}$ values show that PAS is able to predict the target extreme MPM values with an accuracy within 10% for all variations of cases 3 and 4, whereas AS shows deviations of up to 40%. The figure also clearly shows the inherent non-conservative tendency of AS (which is problematic for design load estimation). This trend is not present in the results from PAS; the small deviations observed in the MPM and the distributions can be either conservative or non-conservative. Similar results are observed for $M_{3}$, which highlights that the maximum deviations of the PAS distributions compared to the ground truth are within 18% for all test cases, and those of AS within almost 80%. It should be kept in mind that the present results rely on combining both PAS and AS with linear potential flow indicators. [VS2025] demonstrated that AS, when combined with a coarse-mesh CFD indicator, can yield accurate green water load distributions. However, obtaining indicators from 3D coarse-mesh CFD is far more computationally expensive than using linear potential flow, which highlights the advantage of PAS over AS.
@@
When comparing the extreme values from the validation datasets with those obtained using (P)AS, the respective uncertainties of these values must be considered. The distributions in Figures˜7, 9, 11 and 13 also show the U95% MF-GPR uncertainty for AS, and the U95% sample uncertainty based on 20 draws from the fitted copula model for PAS. These uncertainty measures are not directly comparable, but they show that the sample uncertainty from PAS is small around $P_{\text{exp}}$ for all eight case variations (whereas the MF-GPR uncertainty band from AS is larger around $P_{\text{exp}}$). For case 1, the HF validation sample uncertainty is 0.1 m ($\sim$1% of the true MPM value, see Table˜3 and [VS2025]). The U95% uncertainty of the PAS results is 0.06 m (case 1a) and 0.21 m (case 1b), which corresponds to $\sim$0.6% and $\sim$2% of the true MPM values, respectively. For case 3, the root-mean-square sample errors (RMSE) of the extreme values in the HF validation data are discussed in detail in [VSS2023], showing that these converge to well below 100 kN for the 30-minute MPM in both test conditions. This roughly translates [^2] Assuming a normal distribution, RMSE is equal to standard deviation error, and U95% $\approx 1.96$ RMSE. to an U95% uncertainty around 200 kN ($\sim$7-15% of the true MPM values). The U95% uncertainty of the PAS results is 96 kN (case 3a) and 292 kN (case 3b), which corresponds to $\sim$7% and $\sim$10% of the true MPM values, respectively. The uncertainty of the extreme values in the HF validation data was not evaluated for cases 2 and 4, but it is expected to be between case 1 and 3 for case 2, and comparable to case 3 for case 4. The validation uncertainties reported above for cases 1 and 3 relate to the sample variability over many different realisations of the same wave condition. This does not relate directly to the present validation, as we reproduced the same wave realisation with the LF tools (see Section˜4). We should therefore consider the repetition uncertainty of the experimental results in case 3. We do not have this information directly, but we know from e.g., [Vprads2019] that this uncertainty can be in the same order as the sample variability for wave impacts measured in a seakeeping basin. Based on this, both the experimental and PAS uncertainties are considered to be sufficiently small for all cases to make the validation comparison meaningful.
@@
When accuracy would be the only indicator for the performance of an extreme value prediction method, MCS would always be preferred. The objective of (P)AS is to make accurate predictions in a more efficient way. $M_{1}$ in Figure˜14 shows that both methods significantly reduce the required number of HF samples that need to be evaluated compared to the HF validation data: the number of wave events to evaluate with an HF tool is below 1.7% compared to those required in the MCS for all cases and both AS and PAS. The figure also shows that AS is consistently more efficient than PAS when combined with potential flow indicators. However, [VS2025] demonstrated that using the coarse-mesh CFD indicators required to obtain accurate results for the more complex cases substantially reduce the efficiency of AS. Still, it can be interesting to use AS with potential flow for relatively simple problems such as cases 1 and 2, where it can provide accurate results, and quicker than PAS. PAS is the more reliable method for accurate and efficient predictions across all cases, particularly the complex ones. PAS needs evaluation of less than 0.3% of the HF events compared to MCS for case 1a,b and 2a,b, and less than 1.7% of the HF events compared to MCS for case 3a,b and 4a,b. The exact number of required events depends on the limits set for the stopping criteria. These were defined per case in the present study; the optimal stopping limits for a new case depend on the designer’s requirements, acceptable risks, considered quantiles, and probably requirements by class societies.
@@ E.2 Influence of selected copula model on results
Declaration of generative AI / AI-assisted technology in the manuscript preparation process
During the preparation of this work the authors used ChatGPT in order to rephrase a few individual sentences. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.